If Sally can paint a house in 4 hours, and John can paint the same house in 6 hour, how long will it take for both of them to paint the house together?
How do I set this up to solve it?

Answer by stanbon(57384)

If Sally can paint a house in 4 hours, and John can paint the same house in 6 hour, how long will it take for both of them to paint the house together?
---
Sally rate = 1/4 job/hr
John rate = 1/6 job/hr
--
Together rate: 1/x job/hr
=============================
Equation:
rate + rate + together rate
1/4 + 1/6 = 1/x
Multiply thru by 12x:
3x + 2x = 12
5x = 12
x = 2.4 hr = 2hr 24 min (time to do the job together)
-------------------
Cheers,
Stan H.